now in this equation 9+4(8^2)-30.is the answer 289? or 0? i used 8 to the second power which is 8 times 2 my friend says you have to use square whih is 8 times 8. whos right? 289 or 0? thank you
Answer by stanbon(75887)

9+4(8^2)-30
---
= 9 + 4(64) -30
---
= 9 + 256 - 30
---
= 265 - 30
--------------
= 235
======================
Cheers,
Stan H.
======================